OpenStudy (jessiegonzales):

MEDAL!!!!!!!!!!!!!!!!!!!!!!!!! what is the solution to the equation? 1/x - 6/x^2 =-12

OpenStudy (kropot72):

Is this it? \[\frac{1}{x}-\frac{6}{x ^{2}}=-12\]

OpenStudy (jessiegonzales):

yes

OpenStudy (anonymous):

I think its x=−3/4,2/3 :)

OpenStudy (kropot72):

Okay. First put the left hand side over a common denominator: \[\frac{x-6}{x ^{2}}=-12\] Next cross multiply to get: \[x-6=-12x ^{2}\] Then rearrange to get the quadratic: \[12x ^{2}+x-6=0\] Now you can factorize and solve. Can you do that?

OpenStudy (kropot72):

Is the above method how you got your answer?

OpenStudy (kropot72):

The factorization is: (3x - 2)(4x + 3) = 0 You can now check your solution yourself :)